# An EPTAS for multiprocessor scheduling with rejection under a machine cost constraint

### v1

@@ Abstract
We study the multiprocessor scheduling with rejection problem under a machine cost constraint. In this problem, each job is either rejected with a rejection penalty or; accepted and scheduled on one of the machines for processing. The machine cost is proportional to the total processing time of the jobs scheduled on it. The problem aims to minimize the makespan of accepted jobs plus the total rejection penalty of rejected jobs while the total machine cost does not exceed a given upper bound. We present a simple $2$-approximation algorithm for the problem and we achieve an EPTAS when the number $m$ of machines is a fixed constant.
Keywords: Scheduling within rejection; rejection penalty; green manufacturing; approximation scheme
@@ 1 Introduction
In this paper, we consider the multiprocessor scheduling with rejection problem under a machine cost constraint [12]. Formally, we are given a set of jobs $J=\{J_{1},\ldots,J_{n}\}$ and a set of identical machines $M=\{M_{1},\ldots,M_{m}\}$. A job $J_{j}$ is either rejected with a rejection penalty $e_{j}$ or accepted and scheduled on some machine with a processing time $p_{j}$. Each machine $M_{i}$ is associated with an energy cost rate $r_{i}$ and the machine cost of $M_{i}$ is exactly $r_{i}$ times the maximum completion time of the jobs scheduled on $M_{i}$. The total machine cost is limited and cannot exceed a given upper bound $\hat{U}$. The problem aims to find a solution to minimize the makespan of accepted jobs plus the total rejection penalty of rejected jobs while the machine cost constraint is satisfied. The problem is denoted as $Pm\mid rej,\hat{U},r_{i}\mid C_{\max}(A)+\sum_{J_{j}\in R}e_{j}$ where $m$ is a fixed constant; $r_{i}$ can be non-uniform; $A$ and $R$ are the accepted and rejected jobs, respectively. If $m$ is part of the input, then we use $P$ to replace $Pm$ and if $r_{i}$ is uniform, then we use $r_{i}=r$ to replace $r_{i}$ in the above three-field notation.
The problem is motivated by the applications in the green manufacturing industry [12], which recently attracts much attention because of the growth concerns about environmental challenges. Generally speaking, green manufacturing industry aims to maximize the resource utilities while reduces the environmental damage as much as possible [14]. The machine cost is an important indicator since it may measure the energy cost for the machines, for example, the clean-up cost of pollutants caused by the machines [14]. The authors in [14] initiated the research for scheduling problems in the green manufacturing industry, i.e., $P\mid\hat{U},r_{i}\mid C_{\max}$ and $P\mid\hat{U},r_{i}\mid\sum C_{j}$ where $C_{\max}$ stands for the makespan objective and $\sum C_{j}$ stands for the min-sum objective. Recently, Kong et al. [12] generalized the problem to $P\mid rej,\hat{U},r_{i}=r\mid C_{\max}(A)+\sum_{J_{j}\in R}e_{j}$ by assuming jobs can be either rejected or accepted; but the machine energy cost rate is uniform.
@@
To our knowledge, the only current best $2$-approximation algorithm for $P\mid rej,\hat{U},r_{i}=r\mid C_{\max}(A)+\sum_{J_{j}\in R}e_{j}$ is due to reference [12]. We next review the related works of $P\mid rej,\hat{U},r_{i}=r\mid C_{\max}(A)+\sum_{J_{j}\in R}e_{j}$. Note that when $\hat{U}=\infty$ and each rejection penalty $e_{j}=\infty$, the scheduler is forced to accept all jobs and thus the problem is reduced to the classic multiprocessor scheduling problem $P\mid\mid C_{\max}$ [6]. The problem $P\mid\mid C_{\max}$ is one of the most significant NP-hard problems that has been well studied in the literature. For any fixed $m\geq 2$, $Pm\mid\mid C_{\max}$ is weakly NP-hard [6], for which Sahni presented an FPTAS [15]. Moreover, $P\mid\mid C_{\max}$ has been shown to be strongly NP-hard [11]. Graham [9] designed the first algorithm, called list scheduling, for $P\mid\mid C_{\max}$ in 1966, which always schedules a job on a machine with the least load and achieves $(2-\frac{1}{m})$-approximation. Afterwards, the same author [10] showed that if the jobs are sorted in the non-increasing order of their processing times, the approximation ratio can be improved to $\frac{4}{3}-\frac{1}{3m}$. Finally, Hochbaum and Shmoys [11] resolved this problem by presenting a PTAS.
When $\hat{U}=\infty$, the problem $P\mid rej,U,r_{i}=r\mid C_{\max}(A)+\sum_{J_{j}\in R}e_{j}$ is reduced to the scheduling with rejection problem proposed by Bartal et al. [3]. The authors in [3] presented an FPTAS when $m$ is a fixed constant and a PTAS when $m$ is part of the input. The authors in [12] initiated the research on a general $\hat{U}$ and presented a $2$-approximation algorithm for both a single machine and multiple machines. The interested readers may refer to the survey [16] for more related problems of scheduling with rejection and their corresponding results.
@@ 1.1 Other related works
Note that in a scheduling with rejection problem, the rejection penalty is directly added to the objective if a job is rejected. Many different settings are also considered in the literature. For example, in a scheduling with testing problem, each job $J_{j}$ is given an upper bound $u_{j}$ on the exact processing time $p_{j}$ and a testing time $t_{j}$. A job $J_{j}$ is either untested and processed for $u_{j}$ time or; tested for $t_{j}$ time and processed for $p_{j}$ time on some machine. It is worth mentioning that the testing time $t_{j}$ is added to the processing time of $J_{j}$ instead of the objectives. Dürr et al. [5] first introduced the scheduling with testing problem on a single machine where the testing times are uniform and the goal is to minimize the total completion time. Albers and Eckl [1] extended the model by allowing the testing times to be non-uniform. Moreover, they [2] also pioneered research on minimizing the makespan on multiple machines. Other algorithmic results for the scheduling with testing problems can be found in the references [8, 7].
@@ 1.2 Our results
Our main contribution is an EPTAS for the problem $Pm\mid rej,\hat{U},r_{i}=r\mid C_{\max}(A)+\sum_{J_{j}\in R}e_{j}$, where the number of machines is a fixed constant. Before the EPTAS, we present a new and simpler $2$-approximation algorithm than the one in [12] for $P\mid rej,\hat{U},r_{i}=r\mid C_{\max}(A)+\sum_{J_{j}\in R}e_{j}$ by guessing the maximum processing time of accepted jobs and the maximum rejection penalty of rejected jobs. Our EPTAS first distinguishes the risky jobs whose rejection penalties are relatively big and constructs another instance such that we can easily make the rejection decisions for risky jobs. We next carefully determine the assignments of the ”large” risky jobs that are accepted and present a linear program to deal with the remaining jobs. The jobs with a fractional solution are either rejected or; accepted and appended to the schedule, leading to a small increase on the cost.
The rest of the paper is organized as follows. In Section 2, we give the formal definitions and notation for the problem. In Section 3, we present a new and simple $2$-approximation algorithm and then in Section 4, the EPTAS for $Pm\mid rej,\hat{U},r_{i}=r\mid C_{\max}(A)+\sum_{J_{j}\in R}e_{j}$. Finally, we conclude our paper and introduce some other research problems that can be studied in the future.
@@ 2 Preliminaries
We first present the formal definitions for the scheduling with rejection problem. An instance $I$ of the problem consists of a set of $n$ jobs $J=\{J_{1},\ldots,J_{n}\}$ and a set of $m$ identical machines $M=\{M_{1},\ldots,M_{m}\}$. Each job $J_{j}$ has a known processing time $p_{j}$ and a rejection penalty $e_{j}$. The scheduler needs to make a rejection decision for every job $J_{j}$ and in other words, a job $J_{j}$ is either rejected or accepted. If $J_{j}$ is rejected, then the rejection penalty should be paid; otherwise, $J_{j}$ is accepted and should be scheduled on some machine for processing. The completion time of a machine $M_{i}$, denoted as $t_{i}$, is the maximum completion time of the jobs scheduled on $M_{i}$, that is, the total processing time of the jobs scheduled on $M_{i}$. The makespan of a schedule is the maximum completion time of all machines.
In the green manufacturing industry, the energy cost caused by a machine is usually proportional to its completion time and the total energy cost cannot exceed a given threshold $\hat{U}$. We assume that the energy cost rate for each machine is uniform and is denoted as $r$. Therefore, we have the constraint ”$r\sum_{i=1}^{m}t_{i}\leq\hat{U}$” on the total energy cost [12]. To simplify the notations, we let $U=\hat{U}/r$ and the constraint becomes $\sum_{i=1}^{m}t_{i}\leq U$, which is equivalent to saying that the total processing time of accepted jobs is bounded by $U$. A feasible solution needs to determine the rejection decision for each job and schedule the accepted jobs on the machines while the constraint is satisfied. We define the cost $Z$ of a feasible solution is the makespan of accepted jobs plus the total rejection penalty of rejected jobs. Our goal is to find a feasible solution such that the cost is minimized. If we let $A$ and $R$ denote the accepted and rejected jobs, respectively, then the problem is denoted as $Pm\mid rej,U,r_{i}=r\mid C_{\max}(A)+\sum_{J_{j}\in R}e_{j}$ when the number $m$ of machines is a fixed constant or $P\mid rej,U,r_{i}=r\mid C_{\max}(A)+\sum_{J_{j}\in R}e_{j}$ when $m$ is part of the input.
Given an instance $I$ of $P\mid rej,U,r_{i}=r\mid C_{\max}(A)+\sum_{J_{j}\in R}e_{j}$, we fix an optimal solution of $I$ for discussion and use $A^{*}(I)$ and $R^{*}(I)$ to denote the accepted and rejected job set in the optimal solution, respectively. Moreover, we let $C^{*}_{\max}(I)$ denote the makespan of the optimal solution and $Z^{*}(I)$ denote the optimal cost. Similarly, for a specific solution produced by an algorithm, we let $C_{\max}(I)$ denote the makespan and $Z(I)$ denote the cost of the solution. We drop “$(I)$” from these notations if the instance $I$ is clear from the context.
@@
(List scheduling rule [9]) The List scheduling rule, LS rule for short, always schedules an unassigned job to a machine with the least load.
@@ 3 A new $2$-approximation for $P\mid rej,U,r_{i}=r\mid C_{\max}(A)+\sum_{J_{j}\in R}e_{j}$
We remark that the algorithm in this section can be applied to $P\mid rej,U,r_{i}=r\mid C_{\max}(A)+\sum_{J_{j}\in R}e_{j}$ where the number $m$ of machines is part of the input. Recall that the accepted and rejected job sets are denoted as $A^{*},R^{*}$ in the optimal solution, respectively. Let $p^{*}_{\max}$ denote the maximum processing time of the jobs in $A^{*}$ and $e^{*}_{\max}$ denote the maximum rejection penalty of the jobs in $R^{*}$, respectively. The key idea is to guess the values of $p^{*}_{\max}$ and $e^{*}_{\max}$ so that we can immediately reject the jobs with $p_{j}>p^{*}_{\max}$ and accept the jobs with $e_{j}>e^{*}_{\max}$. The following definition is directly from the fact $p^{*}_{\max}\in\{p_{1},\ldots,p_{n}\}$ and $e^{*}_{\max}\in\{e_{1},\ldots,e_{n}\}$.
@@
A tuple $(p,e)$ is a guess if $p\in\{p_{1},\ldots,p_{n}\}$ and $e\in\{e_{1},\ldots,e_{n}\}$.
It is easy to see that the number of distinct guesses is bounded by $n^{2}$. Note that we do not know $p^{*}_{\max}$ and $e^{*}_{\max}$ and we choose to enumerate all possible guesses. In the following, we use a specific guess $(p_{\ell},e_{k})$ for further discussion. Let
@@
We accept all the jobs in $A_{1}$ and reject all the jobs in $R_{1}$. If the guess is correct, i.e., $(p_{\ell},e_{k})=(p^{*}_{\max},e^{*}_{\max})$, then the jobs in $A_{1}$ are accepted and the jobs in $R_{1}$ with $p_{j}>p_{\ell}$ are rejected in the optimal solution. Therefore, a necessary condition for a correct guess is $U>\sum_{J_{j}\in A_{1}}p_{j}$, which induces the following definition.
@@
If $(p_{\ell},e_{k})$ is not a valid guess, then we skip it and try another guess. Recall that we accept all the jobs in $A_{1}$ and reject all the jobs in $R_{1}$ with $p_{j}>p_{\ell}$, the same as in the optimal solution if the guess is correct. Moreover, we also reject the jobs with $p_{j}>me_{j}$ because the rejection penalty is relatively smaller than the processing time.
For ease of presentation, we let $X=J\setminus(A_{1}\cup R_{1})$. We next present a linear program to determine the rejection decisions for most jobs in $X$. In more detail, let $x_{j}\in\{0,1\}$ denote a binary decision variable for the job $J_{j}\in X$ such that $x_{j}=1$ if and only if $J_{j}$ is rejected by the algorithm. Moreover, let $y_{j}=1-x_{j}$, where $y_{j}$ is also a binary decision variable such that $y_{j}=1$ indicates that $J_{j}$ is accepted. The linear program is presented as follows:
@@
We remark that the objective function is the average processing time on the machines plus the total rejection penalty, which is close to the problem objective $C_{\max}(A)+\sum_{J_{j}\in R}e_{j}$. The first constraint of LP(2) is because the total processing time of the accepted jobs cannot exceed the upper bound $U$ and the jobs in $A_{1}$ have been accepted. Since the guess $(p_{\ell},e_{k})$ is valid, by Definition 3, the linear program has at least one feasible solution, i.e., $x_{j}=1$ for every job $J_{j}\in X$. Note that there are $2|X|\leq 2n$ variables and $|X|+1\leq n+1$ constraints. The linear program is polynomial-time solvable in $n$ because the time complexity of a linear program is polynomial in the number of continuous variables and polynomial in the number of constraints [13].
@@
Note that $\sum_{J_{j}\in X_{1}}p_{j}\leq U-\sum_{J_{j}\in A_{1}}p_{j}$. We reject all the jobs in $X_{2}\cup X_{3}$, in other words, the rejected job set is exactly $R_{1}\cup X_{2}\cup X_{3}$. Finally, we apply the LS rule in Definition 1 to schedule the accepted jobs. The pseudocode of our algorithm is presented as follows.
@@
Input: An instance $J=\{J_{1},\ldots,J_{n}\}$ and $m$ machines;
for Each guess $(p_{\ell},e_{k})$ do
     Define two sets $A_{1},R_{1}$ as in Eq.(1) and $X=J\setminus(A_{1}\cup R_{1})$.
     if $(p_{\ell},e_{k})$ is valid then
         Construct the linear program as in LP(2) and compute a basic solution.
         Partition the jobs in $X$ into three sets $X_{1},X_{2},X_{3}$ as in Eq.(3).
         Reject all the jobs in $X_{2}\cup X_{3}\cup R_{1}$ and apply LS rule to schedule all accepted jobs.
     end if
end for
Return the solution with minimum cost corresponding to the valid guesses.
@@
Consider an assignment of $x_{j}$ for each job $J_{j}\in X$: $y_{j}=1$ if and only if $J_{j}\in A^{*}\cap X$. Once the value of $y_{j}$ is fixed, we set $x_{j}=1-y_{j}$. Note that such an assignment is a feasible solution of LP(2) if the guess is correct. The objective value of the assignment is exactly the right side of Eq.(4), which is an upper bound on the optimal value of LP(2). The proof is done since the left side of Eq.(4) is not greater than the optimal value of LP(2). ∎
@@
The proof is finished. ∎
4 An EPTAS for $Pm\mid rej,U,r_{i}=r\mid C_{\max}(A)+\sum_{J_{j}\in R}e_{j}$
We fix an $10<\epsilon<$ 1 where $\frac{1}{\epsilon}$ is an integer. We assume $m$ is a fixed constant and present an EPTAS for $Pm\mid rej,U,r_{i}=r\mid C_{\max}+\sum_{J_{j}\in R}e_{j}$.

### v2

# A PTAS for multiprocessor scheduling with rejection under a machine cost constraint

@@ Abstract
We study the multiprocessor scheduling with rejection problem under a machine cost constraint. In this problem, each job is either rejected with a rejection penalty or; accepted and scheduled on one of the machines for processing. The machine cost is proportional to the total processing time of the jobs scheduled on it. The problem aims to minimize the makespan of the accepted jobs plus the total rejection penalty of the rejected jobs while the total machine cost does not exceed a given upper bound. We present a simple $2$-approximation algorithm for the problem and we achieve a PTAS even when the number $m$ of machines is part of the input.
Keywords: Scheduling with rejection; rejection penalty; green manufacturing; approximation scheme
@@ 1 Introduction
In this paper, we consider the multiprocessor scheduling with rejection problem under a machine cost constraint [14]. Formally, we are given a set of jobs $J=\{J_{1},\ldots,J_{n}\}$ and a set of identical machines $M=\{M_{1},\ldots,M_{m}\}$. A job $J_{j}$ is either rejected with a rejection penalty $e_{j}$ or accepted and scheduled on some machine with a processing time $p_{j}$. Each machine $M_{i}$ is associated with an energy cost rate $r_{i}$ and the machine cost of $M_{i}$ is exactly $r_{i}$ times the maximum completion time of the jobs scheduled on $M_{i}$. The total machine cost is limited and cannot exceed a given upper bound $\hat{U}$. The problem aims to find a solution to minimize the makespan of the accepted jobs plus the total rejection penalty of the rejected jobs while the machine cost constraint is satisfied. The problem is denoted as $Pm\mid rej,\hat{U},r_{i}\mid C_{\max}(A)+\sum_{J_{j}\in R}e_{j}$ where $m$ is a fixed constant; $r_{i}$ can be non-uniform; $A$ and $R$ are the accepted and rejected jobs, respectively. If $m$ is part of the input, then we use $P$ to replace $Pm$ and if $r_{i}$ is uniform, then we use $r_{i}=r$ to replace $r_{i}$ in the above three-field notation.
The problem is motivated by the applications in green manufacturing industry [14], which recently attracts much attention because of the growth concerns about environmental challenges. Generally speaking, green manufacturing industry aims to maximize the resource utilities while reduces the environmental damage as much as possible [16]. The machine cost is an important indicator since it may measure the energy cost for the machines, for example, the clean-up cost of pollutants caused by the machines [16]. The authors in [16] initiates the research for scheduling problems in green manufacturing industry, i.e., $P\mid\hat{U},r_{i}\mid C_{\max}$ and $P\mid\hat{U},r_{i}\mid\sum C_{j}$ where $C_{\max}$ stands for the makespan objective and $\sum C_{j}$ stands for the min-sum objective. Another classic scheduling problem, referred to as scheduling with rejection [3], assumes the jobs can be rejected or accepted. Such a problem has a natural setting since in many real-life applications, we may have to avoid processing all jobs because of their long processing times even though the rejection penalties must be paid. Recently, Kong et al. [14] combined these two problems and considered a problem $P\mid rej,\hat{U},r_{i}=r\mid C_{\max}(A)+\sum_{J_{j}\in R}e_{j}$ where the machine energy cost rates are uniform.
@@
To our best knowledge, the only and current best $2$-approximation algorithm for $P\mid rej,\hat{U},r_{i}=r\mid C_{\max}(A)+\sum_{J_{j}\in R}e_{j}$ is due to the reference [14]. We next review the related works of $P\mid rej,\hat{U},r_{i}=r\mid C_{\max}(A)+\sum_{J_{j}\in R}e_{j}$. Note that when $\hat{U}$$\infty$ and each rejection penalty $e_{j}=\infty$, the scheduler is forced to accept all jobs and thus the problem is reduced to the classic multiprocessor scheduling problem $P\mid\mid C_{\max}$ [6]. The problem $P\mid\mid C_{\max}$ is one of the most significant NP-hard problems that has been well studied in the literature. For any fixed $m\geq 2$, $Pm\mid\mid C_{\max}$ is weakly NP-hard [6], for which Sahni presented an FPTAS [18]. Moreover, $P\mid\mid C_{\max}$ is proved to be strongly NP-hard [11]. Graham [9] designed the first algorithm, called list scheduling, for $P\mid\mid C_{\max}$ in 1966, which always schedules a job on a machine with the least load and achieves $(2-\frac{1}{m})$-approximation. Afterwards, the same author [10] showed that if the job are sorted in the non-increasing order of their processing times, the approximation ratio can be improved to $\frac{4}{3}-\frac{1}{3m}$. Finally, Hochbaum and Shmoys [11] ended this problem by presenting a PTAS.
When $\hat{U}$$\infty$, the problem $P\mid rej,\hat{U},r_{i}=r\mid C_{\max}(A)+\sum_{J_{j}\in R}e_{j}$ is reduced to the scheduling with rejection problem proposed by Bartal et al. [3]. The authors in [3] presented an FPTAS when $m$ is a fixed constant and a PTAS when $m$ is part of the input. The authors in [14] initiated the research on a general $\hat{U}$ and presented a $2$-approximation algorithm for both a single machine and multiple machines. In [13], the authors presented a very general framework to design an EPTAS for scheduling with rejection and other related problems. We remark that the ”machine activation cost” in [13] is paid for selecting a certain machine type and thus is different from this paper. The interested readers may refer to the survey [19] for more related problems of scheduling with rejection and their corresponding results.
Another problem $P\mid\hat{U},r_{i}\mid C_{\max}$ is also well studied that aims to minimize the makespan but the machine cost is at most $\hat{U}$ with non-uniform machine cost ratios. Li et al. [16] introduced this problem and considered both non-preemptive and preemptive settings for the jobs. They presented an optimal algorithm and a $2$-approximation algorithm for the preemptive and non-preemptive settings, respectively. Later on, Jiang et al. [12] improved the ratio from $2$ to $\frac{\sqrt{33}+1}{4}\approx 1.686$. Li and Ou [17] presented an improved $(\frac{4}{3}+\epsilon)$-approximation algorithm and finally achieved a PTAS by finding a so-called ”monotonic schedule”.
@@ 1.1 Other related works
Note that in a scheduling with rejection problem, the rejection penalty is directly added to the objective if a job is rejected. Many different settings are also considered in the literature. For example, in a scheduling with testing problem, each job $J_{j}$ is given an upper bound $u_{j}$ on the exact processing time $p_{j}$ and a testing time $t_{j}$. A job $J_{j}$ is either untested and processed for $u_{j}$ time or; tested for $t_{j}$ time and processed for $p_{j}$ time on some machine. It is worth mentioning that the testing time $t_{j}$ is added to the processing time of $J_{j}$ instead of the objectives. Dürr et al. [5] first introduced the scheduling with testing problem on a single machine where the testing times are uniform and the goal is to minimize the total completion time. Albers and Eckl [1] extended the model by allowing the testing times to be non-uniform. Moreover, they [2] also pioneered the research for minimizing the makespan on multiple machines. Other algorithmic results for the scheduling with testing problems can be found in the references [8, 7].
@@ 1.2 Our results
Our main contribution is a PTAS for the problem $P\mid rej,\hat{U},r_{i}=r\mid C_{\max}(A)+\sum_{J_{j}\in R}e_{j}$, where the number of machines is part of the input. Before the PTAS, we present a new and simpler $2$-approximation algorithm than the one in [14] for $P\mid rej,\hat{U},r_{i}=r\mid C_{\max}(A)+\sum_{J_{j}\in R}e_{j}$ by guessing the maximum processing time of the accepted jobs and the maximum rejection penalty of the rejected jobs. Our algorithm has a simpler description and avoids the case-by-case analysis in [14]. Our PTAS first distinguishes the risky (safe, respectively) jobs whose rejection penalties are relatively big (small, respectively) and constructs another instance such that we can easily make the rejection decisions for risky jobs. We next partition the safe jobs into a constant number of groups and give a definition of ”rejection types” to describe the number of the accepted jobs in all but one group. Finally, we present a linear program to determine the rejection decisions for almost safe jobs. The jobs with a fractional solution are rejected, leading to a small increase in the cost.
The rest of the paper is organized as follows. In Section 2, we give the formal definitions and notations for the problem. In Section 3, we present a new and simple $2$-approximation algorithm and then in Section 4, the PTAS for $P\mid rej,\hat{U},r_{i}=r\mid C_{\max}(A)+\sum_{J_{j}\in R}e_{j}$. Finally, we conclude our paper and introduce some other research problems that can be studied in the future.
@@ 2 Preliminaries
We first present the formal definitions for the scheduling with rejection problem. An instance $I$ of the problem consists of a set of $n$ jobs $J=\{J_{1},\ldots,J_{n}\}$ and a set of $m$ identical machines $M=\{M_{1},\ldots,M_{m}\}$. Each job $J_{j}$ has a known processing time $p_{j}$ and a rejection penalty $e_{j}$. The scheduler needs to make a rejection decision for every job $J_{j}$ and in other words, a job $J_{j}$ is either rejected or accepted. If $J_{j}$ is rejected, then the rejection penalty should be paid; otherwise, $J_{j}$ is accepted and should be scheduled on some machine for processing. The completion time of a machine $M_{i}$, denoted as $t_{i}$, is the maximum completion time of jobs scheduled on $M_{i}$, that is, the total processing time of jobs scheduled on $M_{i}$. The makespan of a schedule is the maximum completion time of machines.
In green manufacturing industry, the energy cost caused by a machine is usually proportional to its completion time and the total energy cost cannot exceed a given threshold $\hat{U}$. We assume the energy cost rate for each machine is uniform and is denoted as $r$. Therefore, we have the constraint $r\sum_{i=1}^{m}t_{i}\leq\hat{U}$” on the total energy cost [14]. To simplify the notations, we let $U=\hat{U}/r$ and the constraint becomes $\sum_{i=1}^{m}t_{i}\leq U$, which is equivalent to saying that the total processing time of the accepted jobs is bounded by $U$. A feasible solution needs to determine the rejection decision for each job and schedule the accepted jobs on the machines while the constraint is satisfied. We define the cost $Z$ of a feasible solution is the makespan of the accepted jobs plus the total rejection penalty of the rejected jobs. Our goal is to find a feasible solution such that the cost is minimized. If we let $A$ and $R$ denote the accepted and rejected jobs, respectively, then the problem is denoted as $P\mid rej,U,r_{i}=r\mid C_{\max}(A)+\sum_{J_{j}\in R}e_{j}$ when the number $m$ of machines is part of the input.
Given an instance $I$ of $P\mid rej,U,r_{i}=r\mid C_{\max}(A)+\sum_{J_{j}\in R}e_{j}$, we fix an optimal solution of $I$ for discussion and use $A^{*}(I)$ and $R^{*}(I)$ to denote the accepted and rejected job set in the optimal solution, respectively. Moreover, we let $C^{*}_{\max}(I)$ denote the makespan of the optimal solution and $Z^{*}(I)$ denote the optimal cost. Similarly, for a specific solution produced by an algorithm, we let $C_{\max}(I)$ denote the makespan and $Z(I)$ denote the cost of the solution. We drop “(I)(I)” from these notations if the instance $I$ is clear from the context.
@@
(List scheduling rule [9]) The List scheduling rule, LS rule for short, always schedule an unassigned job to a machine with the least load.
@@ 3 A new $2$-approximation for $P\mid rej,U,r_{i}=r\mid C_{\max}(A)+\sum_{J_{j}\in R}e_{j}$
Recall that the accepted and rejected job sets are denoted as $A^{*},R^{*}$ in the optimal solution, respectively. Let $p^{*}_{\max}$ denote the maximum processing time of the jobs in $A^{*}$ and $e^{*}_{\max}$ denote the maximum rejection penalty of the jobs in $R^{*}$, respectively. The key idea is to guess the values of $p^{*}_{\max}$ and $e^{*}_{\max}$ so that we can immediately reject the jobs with $p_{j}>p^{*}_{\max}$ and accept the jobs with $e_{j}>e^{*}_{\max}$. The following definition is directly from the fact $p^{*}_{\max}\in\{0,p_{1},\ldots,p_{n}\}$ and $e^{*}_{\max}\in\{0,e_{1},\ldots,e_{n}\}$ where $p^{*}_{\max}=0$ ($e^{*}_{\max}$, respectively) indicates that all jobs are rejected (accepted, respectively).
@@
A tuple $(p,e)$ is a guess if $p\in\{0,p_{1},\ldots,p_{n}\}$ and $e\in\{0,e_{1},\ldots,e_{n}\}$.
It is easy to see that the number of distinct guesses is bounded by $O(n^{2})$. Note that we do not know $p^{*}_{\max}$ and $e^{*}_{\max}$ and we choose to enumerate all possible guesses. In the following, we use one guess $(p_{\ell},e_{k})$ for further discussion. Let
@@
We accept all jobs in $A_{1}$ and reject all jobs in $R_{1}$. If the guess is correct, i.e., $(p_{\ell},e_{k})=(p^{*}_{\max},e^{*}_{\max})$, then the jobs in $A_{1}$ are accepted and the jobs in $R_{1}$ with $p_{j}>p_{\ell}$ are rejected in the optimal solution. Therefore, a necessary condition for a correct guess is $U>\sum_{J_{j}\in A_{1}}p_{j}$, which induces the following definition.
@@
If $(p_{\ell},e_{k})$ is not a valid guess, then we skip it and try another guess. Recall that we accept all jobs in $A_{1}$ and all jobs in $R_{1}$ with $p_{j}>p_{\ell}$, the same as in the optimal solution if the guess is correct. Moreover, we also reject the jobs with $p_{j}>me_{j}$ because the rejection penalty is relatively smaller than the processing time.
For ease of presentation, we let $X=J\setminus(A_{1}\cup R_{1})$. We next present a linear program to determine the rejection decisions for most jobs in $X$. In more details, let $x_{j}\in\{0,1\}$ denote a binary decision variable for the job $J_{j}\in X$ such that $x_{j}=1$ if and only if $J_{j}$ is rejected by the algorithm. Moreover, let $y_{j}=1-x_{j}$, where $y_{j}$ is also a binary decision variable such that $y_{j}=1$ indicates $J_{j}$ is accepted. We relax the binary variables $x_{j},y_{j}$ to be continuous and the linear program is presented as follows:
@@
We remark that the objective function is the average processing time on the machines plus the total rejection penalty, which is close to the problem objective $C_{\max}(A)+\sum_{J_{j}\in R}e_{j}$. The first constraint of LP(2) is because the total processing time of the accepted jobs cannot exceed the upper bound $U$ and the jobs in $A_{1}$ have been accepted. Since the guess $(p_{\ell},e_{k})$ is valid, by Definition 3, the linear program has at least one feasible solution, i.e., $x_{j}=0$ for every job $J_{j}\in X$. Note that there are $2|X|\leq 2n$ variables and $|X|+1\leq n+1$ constraints. The linear program is polynomial-time solvable in $n$ because the time complexity of a linear program is polynomial in the number of continuous variables and polynomial in the number of constraints [15]. Typically, LP(2) can be solved in $O(n^{3.5}L)$ time using interior point methods where $L$ is the total size of the instance $I$.
@@
Note that $\sum_{J_{j}\in X_{1}}p_{j}\leq U-\sum_{J_{j}\in A_{1}}p_{j}$. We reject all the jobs in $X_{2}\cup X_{3}$, in other words, the rejected job set is exactly $R_{1}\cup X_{2}\cup X_{3}$. Finally we apply LS rule in Definition 1 to schedule the accepted jobs. The pseudo-code of our algorithm is presented in Algorithm 1.
@@
1:Input: An instance $J=\{J_{1},\ldots,J_{n}\}$ and $m$ machines;
2:for Each guess $(p_{\ell},e_{k})$ do
3:  Define two sets $A_{1},R_{1}$ as in Eq.(1) and $X=J\setminus(A_{1}\cup R_{1})$.
4:  if $(p_{\ell},e_{k})$ is valid then
5:   Construct the linear program as in LP(2) and compute a basic solution.
6:   Partition the jobs in $X$ into three sets $X_{1},X_{2},X_{3}$ as in Eq.(3).
7:   Reject all the jobs in $X_{2}\cup X_{3}\cup R_{1}$ and apply LS rule to schedule all the accepted jobs.
8:  end if
9:end for
10:Return the solution with minimum cost corresponding to the valid guesses.
@@
Consider an assignment of $x_{j}$ for each job $J_{j}\in X$: $x_{j}=1$ if and only if $J_{j}\in A^{*}\cap X$. Once the value of $x_{j}$ is fixed, we set $y_{j}=1-x_{j}$. Note that such an assignment is a feasible solution of LP(2) if the guess is correct. The objective value of the assignment is exactly the right side of Eq.(4), which is an upper bound on the optimal value of LP(2). The proof is done since the left side of Eq.(4) is not greater than the optimal value of LP(2). ∎
We are ready to show our algorithm is indeed $2$-approximation.
@@
The proof is finished.
4 A PTAS for $P\mid rej,U,r_{i}=r\mid C_{\max}(A)+\sum_{J_{j}\in R}e_{j}$
Let $0<\epsilon<1$ is a given constant and for simplicity, we assume $\frac{1}{\epsilon}$ is an integer. We consider the case where $m$ is part of the input and present a PTAS for $P\mid rej,U,r_{i}=r\mid C_{\max}+\sum_{J_{j}\in R}e_{j}$.